# Data Quality Matters: Suicide Intention Detection on Social Media Posts Using a RoBERTa-CNN Model

### v1

@@ Abstract
Suicide remains a global health concern for the field of health, which urgently needs innovative approaches for early detection and intervention. In this paper, we focus on identifying suicidal intentions in SuicideWatch Reddit posts and present a novel approach to suicide detection using the cutting-edge RoBERTa-CNN model, a variant of RoBERTa (Robustly optimized BERT approach). RoBERTa is used for various Natural Language Processing (NLP) tasks, including text classification and sentiment analysis. The effectiveness of the RoBERTa lies in its ability to capture textual information and form semantic relationships within texts. By adding the Convolution Neural Network (CNN) layer to the original model, the RoBERTa enhances its ability to capture important patterns from heavy datasets. To evaluate the RoBERTa-CNN, we experimented on the Suicide and Depression Detection dataset and obtained solid results. For example, RoBERTa-CNN achieves 98% mean accuracy with the standard deviation (STD) of 0.0009. It also reaches over 97.5% mean AUC value with an STD of 0.0013. In the meanwhile, RoBERTa-CNN outperforms competitive methods, demonstrating the robustness and ability to capture nuanced linguistic patterns for suicidal intentions. Therefore, RoBERTa-CNN can detect suicide intention on text data very well.
@@
Suicide Intention Detection, NLP, RoBERTa, CNN, Data Quality
@@ I Introduction
Suicide is one of the leading causes of mortality worldwide with approximately 1.53 million cases in 2020 [1]. The World Health Organization (WHO) estimated that approximately 703,000 people committed suicide worldwide in 2022 [2]. At least 1 out of 31 individuals with suicide ideation attempted to end this life [3]. Thus, suicide is an urgent public health concern worldwide, which requires early intervention [4, 5].
@@
Linguistic symptoms like pessimistic note writing and remarks are common signs of committing suicide [4]. Therefore, linguistic symptoms can benefit suicide intention detection (SID) for early intervention. Then, even desperate people, nowadays, are inclined to anonymously express themselves truly and freely online instead of in the real world. Hence, it is more practical to research online content. At the same time, artificial intelligence (AI) and deep machine learning (ML) already started a way for psychology research [6, 7] like SID [8]. Given the importance of SID and the power of AI/ML, this paper presents our recent research on using natural language processing (NLP) models to research SID by analyzing linguistic symptoms from social media posts and comments (Reddit).
@@
Specifically, social media has an abundant linguistic corpus and provides the source for many NLP research [9, 10]. In this paper, we select Suicide and Depression Detection (SDD) dataset [^1] https://www..com/datasets/nikhileswarkomati/suicide-watch. The developer collected this dataset by scrapping the posts and comments from the SuicideWatch and Depression subreddits of the Reddit Platform [^2] https://www.reddit.com/r/SuicideWatch/.
@@
Currently, it is widely accepted that transformer-based NLP models are capable of tackling various medical tasks like cardiology and neurology research  [11] and mental health tasks [12]. To capture the early signs of suicide ideation from linguistic data, we created a RoBERTa-CNN  [13, 4, 14] model to conduct the study. RoBERTa-CNN, as a variant of BERT (Bidirectional Encoder Representations from Transformers), contains the RoBERTa backbone and convolutional head. The attention mechanism enables RoBERTa to comprehend complex contextual information and capture fine linguistic patterns well [15], while the convolutional head strengthens to obtain linguistic patterns and extract speech-related information in order to make the model more robust and generalizable [16, 9]. RoBERTa-CNN not only improves the performance but also overcomes the constraints of the original RoBERTa model [15]. Specifically, the RoBERTa eliminates the next sentence prediction and uses a larger dataset for the pre-trained model, which improves contextual awareness over typical models [15]. Compared to the BERT and RoBERTa models, RoBERTa-CNN has the ability to handle long sequences more effectively [17, 15] due to its convolutional neural network (CNN). Hence, RoBERTa-CNN is well-suited for SID.
@@
Our experimental result also verifies this conclusion. RoBERTa-CNN received a high mean test accuracy of 98%, which outperforms other text-based SID studies (see Table III). This good performance demonstrates that RoBERTa-CNN can predict SID with high accuracy as it is robust under different contexts and expressions.
By the way, this research also states that data quality matters. RoBERTa-CNN is good, but far from a panacea. It is the high data quality contributing to the excellent performance. Good data quality is still the core for AI models to get good predictions.
@@
  • The proposed automatic SID system achieved higher experimental results (98% accuracy) compared to the state-of-the-art methods on the same dataset.
@@
  • We utilized Generative Pre-trained Transformer (GPT) API to clean up the dataset as data quality matters a lot in training the RoBERTa-CNN model.
The remainder of the paper is organized as follows. Section 2 reviews the related works on SID. Section 3 introduces the proposed methodology. Section 4 presents the experiments and the ablation study. Section 5 covers the discussion and future work direction. Finally, Section 6 concludes the paper.
@@ II Related Works
Currently, many researchers turned to BERT-based models [18] for SID [19], which capture semantics and contextual features well [20]. RoBERTa is one of the BERT family, it applies dynamic masking to embed text instead of the static embedding used in the typical BERT. RoBERTa allows the large training batch size to avoid logic errors. Unlike the BERT model, RoBERTa reduces the warming-up steps and increases the learning rate [21, 15], which makes RoBERTa simply and efficient [15].
Then, RoBERTa works well in various NLP tasks such as text summarization [22, 23] and text analysis [24]. Particularly, Mishra et al. [22] applied RoBERTa and achieved 86% accuracy for the factual correctness of textual summarization. Mohamed et al. [23] reported 76% accuracy for text analysis by using the RoBERTa.
Then, ALBERT, XLNet, and DistillBERT, the other BERT variants, also had a robust performance on NLP tasks [18, 25, 26, 27]. However, compared with the RoBERTa, the ALBERT model is time-consuming [28], the XLNet is slower too [15], and the DistillBERT loses information due to the reduction of model size [29, 18].
In addition, many works focused on analyzing keywords instead of the entire sentence [30], which causes information loss. Hence, our study implements RoBERTa as the backbone to analyze the entire sentence for SID.
@@ III Methodology
This paper applies the RoBERTa-CNN model, which involves the RoBERTa backbone and a CNN head. The RoBERTa studies textual features from the dataset. The CNN classifier processes the extracted features and enhances the prediction performance. Sections III-A to III-C illustrate the details of RoBERTa. Section III-D describes the loss function.
@@ III-A RoBERTa Embedding
We applied RoBERTa embedding to tokenize input comments (individuals’ posts on Reddit), $X=[x_{1},x_{2},\dotsm,x_{n}]$, where $n$ is the number of sentences, $x_{i}$ has the size of [Max_Len], $i\in[1,n]$, Max_Len is the token length of $x_{i}$. The shape of $X$ is [n, Max_Len].
Compared with the context-independent word embedding approaches such as GloVE, RoBERTa embedding belongs to context-dependent word embedding [15], which considers the order of words within the context and learns the context-specific nuances [4, 15]. Thereby, $X$ contains many dynamic and robust representations. Moreover, the RoBERTa embedding model is pre-trained on large text corpora.
@@ III-B RoBERTa Backbone
Fig. 1 shows the RoBERTa backbone, which takes $X$ as the input and passes $X$ into the transformer encoder in which the core is the attention mechanism (1).
Given a layer $i$, the computational processes are (2) and (3).
@@
where $L_{i-1}$ is the output of layer $i-1$ and the input of this layer $i$, MHSA is multi-head self-attention module, LN represents layer normalization, $\textbf{Attn}_{out}$ is the output of the attention module. Then, FF is the feed-forward module, and $L_{i}$ is the output of layer $i$. Its shape is [n, Max_Len, 768], where 768 is the hidden dimension within the attention mechanism.
While analyzing text data, the attention mechanism focuses more on suicidal intention-related words of the input sequence, which benefits the RoBERTa-CNN in identifying imperceptible but valuable linguistic signs well  [4, 6, 15].
@@ III-C CNN head
Fig. 1 shows that the CNN head contains one Conv layer, one ReLU layer, one 1D-max pooling (Max_Pool1D) layer, and one fully-connected (FC) layer, which can be summarized as (4) and (5). The CNN head takes $L_{N}$ as the input, where $L_{N}$ is the output of the backbone after swapping the last two dimensions, and $N$ represents the number of attention layers. The size of $L_{N}$ is [n, 768, Max_Len]. The CNN head outputs $\hat{y}$ with the shape of [n, 2].
@@
where the kernel size and channel number of the Conv layer are 2 and 100. After doing max pooling, the size of CNN_Out is [n, 100, 1]. Squeeze reduces CNN_Out from a 3D to a 2D tensor with the shape of [n, 100]. Then, the FC layer deducts dimensions from 100 to 2. The Conv layer enables the RoBERTa-CNN to pay attention to the local features and patterns that are vital to the classification.
III-D Loss Function
We utilize the Binary Cross-Entropy loss (BCE (6)). It measures the cost between ground truth value $y$ and the prediction $\hat{y}$.
$\text{BCE}(y,\hat{y})=\sum^{n}_{i=1}[\hat{y}_{i}\cdot\log y_{i}+(1-\hat{y}_{i})\cdot\log(1-y_{i})]$ (6)
@@ IV Experiment
This section first introduces the SDD dataset and experiment settings. Then, it presents and analyzes the experimental results and ablation study.
@@ IV-A Dataset
The Suicide and Depression Detection (SDD) dataset consists of the online posts and comments from the SuicideWatch and Depression channel, the subreddits of the Reddit Platform. The time range is from December 2018 to January 2021. Table I shows some data samples. After preprocessing, it reminds 110,040 out of 232,074 posts with two categories (Suicide and non-Suicide).
@@ 3. Why is everyone talking about us? I’m genuinely confused what exactly did we do?
Data processing: Three data cleaning steps are: 1) remove all noise symbols (accents, hyperlinks, and emoticons); 2) remove all non-English characters; and 3) delete all sentence fragments. Since the research involves sentence analysis, normal standard preprocessing such as converting all text to lowercase, removing punctuation and accents, stripping white space, and stop-words were all avoided. Therefore, we manually cleaned 50,000 samples and applied GPT API to clean 60,040 other samples. There are 35,270 suicide samples and 74,700 non-suicide ones.
@@
Then, following the pattern of conductive learning, we randomly and stratifiedly split the dataset into non-overlapped three subsets, train, validation, and test sets with the ratio of 8:1:1 [31]. Table II has the distribution.
@@ IV-B Implementation detail
While training, we augmented the text first with normalization, retrieval of synonyms, and synonym replacement. The batch size is 200 and the optimizer is the Adam optimizer with the initialized learning rate of 1e-6. Conductive learning tests if RoEBRTa-CNN can detect unseen data well. Then, we randomly trained the RoBERTa-CNN (over 35 epochs) on the SID dataset using an NVIDIA RTX 3090 GPU five times and reported the mean and standard deviation (STD) values.
@@ IV-C Evaluation metrics
The used metrics are Accuracy, Precision, Recall, F1 Score, and AUC (Area Under the Receiver Operating Characteristic Curve). The higher these metric values are, the better and more robust the RoBERTa-CNN is.
@@ IV-D Experiment by RoBERTa-CNN
To evaluate RoBERTa-CNN on SID, we experimented with the SDD dataset. Table III shows that RoBERTa-CNN achieved 98.00% mean accuracy with an STD of 0.09%, which is 3% and 6.5% higher than the accuracy on the CNN-BiLSTM and XGBoost models, respectively. Then, the overall highest Precision, Recall, F1 Score, and AUC also indicate that RoBERTaCNN is very accurate and robust at detecting both suicide and non-suicide intention. In addition, the conductive learning setting makes the experimental result more objective and convincing as well. Hence, RoBERTa-CNN is capable of predicting SID, and it surpasses CNN-BiLSTM and XGBoost [32] on the SDD dataset.
@@ IV-E Ablation study
Each sample’s token length, denoted as Max_Len, significantly affects the performance of RoBERTa-CNN. By setting Max_Len as $[64,128,256]$, we did an ablation study to find optimized Max_Len. Table IV illustrates the results from the validation and test sets.
TABLE IV: The experimental results of different Max_Len on validation and test sets. The value format is Mean±$\pm$±STD.
Validation Set
Max_Len Accuracy (%) Precision(%) Recall (%) F1 Score (%) AUC(%)
64 98.09$\pm$0.06 97.31$\pm$0.42 96.69$\pm$0.40 96.99$\pm$0.08 97.72$\pm$0.11
128 98.09$\pm$0.09 96.69$\pm$0.39 97.32$\pm$0.19 97.00$\pm$0.14 97.88$\pm$0.06
256 98.04$\pm$0.04 96.94$\pm$0.42 96.92$\pm$0.37 96.93$\pm$0.06 97.74$\pm$0.09
Test Set
@@
Table IV shows that, on the validation set, RoBERTa-CNN with the 128 Max_Len surpasses that with 64 and 256 ones. In detail, the 128 Max_Len helps the model get the four best metrics, which are accuracy ((98.09$\pm$0.09)%), Recall, F1 Score, and AUC. In addition, the model with the 256 Max_Len has the second-best Precision, Recall, and AUC. Overall, 256 is the second suitable Max_Len value on the validation set.
On the test set, Table IV demonstrates that the 256 Max_Len leads to the best performance with the top 1 accuracy of (98.00$\pm$0.09)and the top 1 F1 Score of (96.81$\pm$0.14)%. Furthermore, 256 Max_Len achieves comparable Precision, Recall, and AUC, which are very close to the highest values. For example, the mean Precision of 256 Max_Len (96.98%) is merely 0.01% lower than that of 64 one, and the STD of 256 Max_Len (0.0022) is 0.0023 lower than that of 64 one (0.0045). This means that 256 Max_Len contributes to a more robust Precision value among 5 experiments than 64 Max_Len. In general, 256 Max_Len suits more for the test set.
Due to the conductive learning, we weigh the performance of the test set more than that of the validation one. Therefore, we set the Max_Len as 256 in the RoBERTa-CNN. In the meanwhile, overall good results of the ablation study show that RoBERTa-CNN is robust on the SDD dataset.

### v2

# Data Quality Matters: Suicide Intention Detection on Social Media Posts Using RoBERTa-CNN

@@ Abstract
Suicide remains a pressing global health concern, necessitating innovative approaches for early detection and intervention. This paper focuses on identifying suicidal intentions in posts from the SuicideWatch subreddit by proposing a novel deep-learning approach that utilizes the state-of-the-art RoBERTa-CNN model. The robustly Optimized BERT Pretraining Approach (RoBERTa) excels at capturing textual nuances and forming semantic relationships within the text. The remaining Convolutional Neural Network (CNN) head enhances RoBERTa’s capacity to discern critical patterns from extensive datasets. To evaluate RoBERTa-CNN, we conducted experiments on the Suicide and Depression Detection dataset, yielding promising results. For instance, RoBERTa-CNN achieves a mean accuracy of 98% with a standard deviation (STD) of 0.0009. Additionally, we found that data quality significantly impacts the training of a robust model. To improve data quality, we removed noise from the text data while preserving its contextual content through either manually cleaning or utilizing the OpenAI API.
@@
Suicide Intention Detection, NLP, RoBERTa, CNN, Data Quality, OpenAI API
@@ I Introduction
Suicide is one of the leading causes of mortality worldwide with approximately 1.53 million cases in 2020 [1]. The World Health Organization (WHO) estimated that approximately 703,000 people committed suicide worldwide in 2022 [2]. At least 1 out of 31 individuals with suicide ideation attempted to end their life [3]. Thus, suicide is an urgent public health concern worldwide, which requires early intervention [4].
@@
Linguistic symptoms such as pessimistic note writing and remarks are common signs of committing suicide ideation [4], which, thus, benefits suicide intention detection (SID) for early intervention. Nowadays, people are inclined to anonymously express themselves freely online instead of in the real world. Hence, it is more practical to research online content. At the same time, deep learning (DL) already paved the way for psychology research like SID [5]. Given the importance of SID and the power of DL, this paper presents our recent study on using natural language processing (NLP) models to research SID by analyzing linguistic symptoms from social media posts and comments (Reddit).
@@
Specifically, social media has an abundant linguistic corpus for NLP research [6]. This paper utilized the Suicide and Depression Detection (SDD) dataset [^1] https://www..com/datasets/nikhileswarkomati/suicide-watch. The developer collected SDD by scrapping the posts and comments from the SuicideWatch and Depression subreddits of the Reddit Platform [^2] https://www.reddit.com/r/SuicideWatch/.
@@
Currently, transformer-based NLP models are good at tackling various medical tasks (cardiology and neurology research [7]) and mental health tasks [8]. To capture the early signs of suicide ideation from linguistic data, we created a RoBERTa-CNN [4] model to conduct the study. RoBERTa-CNN, as a variant of BERT (Bidirectional Encoder Representations from Transformers), contains RoBERTa backbone and CNN head. The attention mechanism enables RoBERTa to comprehend complex contextual information and capture sophisticated linguistic patterns well [9], while the CNN head strengthens the power to obtain linguistic patterns and extract speech-related features, making the model more robust and generalizable [10]. RoBERTa-CNN improves the performance and overcomes the constraints of the original RoBERTa model [9]. Specifically, RoBERTa eliminates the next sentence prediction and uses a larger dataset for the pre-trained model, which improves contextual awareness over typical models [9]. Compared to BERT and RoBERTa, RoBERTa-CNN handles long sequences more effectively [11, 9] due to its CNN head. Hence, RoBERTa-CNN is well-suited for SID.
@@
RoBERTa-CNN achieved a high mean accuracy of 98% in our experiments, which outperforms other text-based SID studies (see Table III). Furthermore, this solid performance demonstrates that RoBERTa-CNN predicts SID reliably as it is robust under different contexts and expressions.
This research also supports the significance of data quality. Our data cleaning process eliminates noisy information, thus enhancing data quality and contributing to superior performance. This underscores the essential role of providing high-quality data for training AI models to achieve high accuracy.
@@
  • The proposed automatic SID system achieved a mean accuracy of 98%, higher than that in the state-of-the-art methods on the same dataset.
@@
  • We used OpenAI API to clean up the dataset due to the key of data quality in training RoBERTa-CNN.
The remainder of this paper is organized as follows. Section II reviews the related works on SID. Section III introduces the model. Section IV presents the experiments and the ablation study. Sections V and VI cover the discussion and conclusion.
@@ II Related Works
Currently, BERT-based models [12] are good for SID [13], which capture semantics and contextual features well [14]. RoBERTa, BERT’s variant, applies dynamic masking to embed text instead of the static embedding used in the typical BERT. RoBERTa utilizes the large training batch size to avoid logic errors. Unlike BERT, RoBERTa reduces the warming-up steps and increases the learning rate [15, 9], which makes RoBERTa efficient [9].
RoBERTa succeeds in various NLP tasks such as text summarization [16, 17] and text analysis [18]. Specifically, Mishra et al. [16] applied RoBERTa and achieved 86% accuracy for the factual correctness of textual summarization. Mohamed et al. [17] reported 76% accuracy for text analysis by using RoBERTa.
Other BERT variants perform robustly on NLP tasks [12, 19, 20, 21] but have drawbacks. Compared with RoBERTa, ALBERT is time-consuming [22], XLNet is slower too [9], and DistillBERT loses information due to the reduction of model size [12].
In addition, much work has focused on analyzing keywords instead of the entire sentence [23], which causes information loss. Hence, this study uses RoBERTa-Base as the backbone to do sentence analysis.
@@ III Methodology
This paper applies RoBERTa-CNN, which involves a RoBERTa backbone and a CNN head. RoBERTa extracts textual features from the dataset (Sections III-A and III-B). The CNN classifier processes the extracted features and enhances the prediction performance (Section III-C).
@@ III-A RoBERTa Embedding
The RoBERTa embedding tokenizes input comments (individuals’ posts on Reddit), $X=[x_{1},x_{2},\dotsm,x_{n}]$, where $n$ is the number of sentences, $x_{i}$ has the size of [Max_Len], $i\in[1,n]$, Max_Len is the token length of $x_{i}$. The shape of $X$ is [n, Max_Len].
The RoBERTa embedding is a context-dependent word embedding [9], which considers word order within the context and learns context-specific nuances [4, 9]. Thereby, $X$ contains many dynamic and robust representations. Moreover, the RoBERTa embedding is pre-trained on large text corpora.
@@ III-B RoBERTa Backbone
Fig. 1 shows that RoBERTa passes the input $X$ into the transformer encoder where the core is the attention mechanism (1). Given a layer $i$, the computing processes are (2) and (3).
@@
where $L_{i-1}$ is the output of layer $i-1$ and the input of this layer $i$, MHA is multi-head self-attention module, LN represents layer normalization, $\text{Attn}_{out}$ is the output of the attention module. FF is the feed-forward module, and $L_{i}$ is the output of layer $i$. Its shape is [n, Max_Len, 768], where 768 is the hidden dimension within the attention mechanism.
While analyzing text data, the attention mechanism focuses more on suicidal intention-related words, which benefits RoBERTa-CNN in identifying imperceptible but valuable linguistic signs well  [4, 9].
@@ III-C CNN head
The CNN head contains one Conv layer, one ReLU layer, one 1D-max pooling (Max_Pool1D) layer, and one fully-connected (FC) layer (See Fig. 1), which is summarized as (4) and (5). The CNN head’s input is RoBERTa’s output $L_{N}$ after swapping the last two dimensions. $N$ is the number of attention layers. The size of $L_{N}$ is [n, 768, Max_Len]. The CNN head outputs $\hat{y}$ with the shape of [n, 2].
@@
The kernel size and channel number of the Conv layer are 2 and 100. The size of CNN_Out is [n, 100, 1]. Squeeze function reduces CNN_Out from a 3D to a 2D tensor with the size of [n, 100]. Then, the FC layer deducts dimensions from 100 to 2. The Conv layer enables RoBERTa-CNN to pay attention to the local features and patterns that are key for the classification.
@@ IV Experiment
This section first introduces the SDD dataset and the experiment settings. Then, it presents and analyzes the experimental results and ablation study.
@@ IV-A Dataset
The Suicide and Depression Detection (SDD) dataset consists of the online posts and comments from the SuicideWatch and Depression channel, subreddits of the Reddit Platform. The time range is from December 2018 to January 2021. Table I shows some data samples. We randomly picked 110,040 out of 232,074 posts with two categories (Suicide and non-Suicide).
@@ 3. Why is everyone talking about us? I’m genuinely confused what exactly did we do?
Data processing: Three data cleaning steps are: 1) remove all noise symbols (accents, hyperlinks, and emoticons); 2) remove all non-English characters; and 3) delete all sentence fragments. Since the research involves sentence analysis, normal standard preprocessing was avoided, like converting all text to lowercase, removing punctuation and accents, stripping white space, and stop-words. We manually cleaned 50,000 samples and applied OpenAI API to clean 60,040 others. There are 35,270 suicide samples and 74,770 non-suicide ones.
@@
Then, following the pattern of conductive learning, we randomly and stratifiedly split the dataset into non-overlapped subsets, train, validation, and test sets, with the ratio of 8:1:1 [24]. Table II presents the distribution.
@@ IV-B Implementation detail
For training, we first augmented the text with normalization, retrieval of synonyms, and synonym replacement. The batch size was 200. We initialized the learning rate to 1e-6 for the Adam optimizer. Conductive learning tests if RoEBRTa-CNN can detect unseen data well. Binary Cross-entropy loss controls the training. Then, we randomly trained RoBERTa-CNN (over 35 epochs) on the SID dataset using an NVIDIA RTX 3090 GPU five times and reported the mean and standard deviation (STD) values.
@@ IV-C Evaluation metrics
The used metrics are Accuracy, Precision, Recall, F1 Score, and AUC (Area Under the Receiver Operating Characteristic Curve). Higher metric values mean better RoBERTa-CNN.
@@ IV-D Experiment by RoBERTa-CNN
The SDD dataset is to evaluate RoBERTa-CNN on SID. Table III implies that RoBERTa-CNN achieved 98.00% mean accuracy with an STD of 0.09%, which is 3%, 6.5%, and 1.8% higher than the accuracy on CNN-BiLSTM, XGBoost, and EFS + pBGSK + SVM, respectively. Then, the overall highest Precision, Recall, F1 Score, and AUC also indicate that RoBERTa-CNN is very accurate and robust at detecting both suicide and non-suicide intention. In addition, the conductive learning setting makes the experimental result more objective and convincing. Hence, RoBERTa-CNN predicts SID well and surpasses CNN-BiLSTM, XGBoost, and EFS + pBGSK + SVM on the SDD dataset.
@@ IV-E Ablation study
IV-E1 The Effectiveness of Max_Len
Each sample’s token length $Max\_Len$ significantly affects the model performance. By setting Max_Len as $[64,128,256]$, we did an ablation study to find optimized Max_Len.
TABLE IV: The ablation study results. The upper table is to tune Max_Len, and the bottom one is to compare with the origin RoBERTa. The value format is Mean±$\pm$±STD.
@@ IV-E1 The Effectiveness of Max_Len
Classifier Accuracy (%) Precision(%) Recall (%) F1 Score (%) AUC(%)
RoBERTa 97.95$\pm$0.06 97.03$\pm$0.39 96.70$\pm$0.39 96.86$\pm$0.05 97.66$\pm$0.11
RoBERTa-CNN 97.98$\pm$0.06 96.74$\pm$0.41 96.87$\pm$0.28 96.80$\pm$0.09 97.68$\pm$0.06
Table IV reveals that setting Max_Len = 256 leads to the best performance with a Top-1 accuracy of (98.00$\pm$0.09)and a Top-1 F1 Score of (96.81$\pm$0.14)%. Furthermore, setting Max_Len = 256 achieves comparable Precision, Recall, and AUC, which are very close to the highest values. For example, the mean Precision (96.98%) when Max_Len = 256 is merely 0.01% lower than that of Max_Len = 64, and the STD of Max_Len = 256 (0.22%) is 0.23% lower than that of Max_Len = 64 (0.45%). Hence, Max_Len = 256 contributes to a more robust Precision value among five experiments than Max_Len = 64. In general, choosing Max_Len = 256 suits the model performance more.
IV-E2 The Effectiveness of CNN Classifier
This experiment tests whether or not the CNN classifier strengthens RoBERTa. To speed up the experiment, we set Max_Len to 128.
Table IV implies that RoBERTa-CNN (97.98%) reaches better mean accuracy than RoBERTa (97.95%). RoBERTa-CNN’s mean AUC and mean Recall are higher than RoBERTa’s, which indicates that the CNN classifier detects the positive and negative samples better than the classifier used by RoBERTa. Moreover, RoBERTa-CNN’s solid Precision and F1 Score are comparable to RoBERTa’s. Hence, RoBERTa-CNN, in comparison, is better than RoBERTa on the SDD dataset.
Overall good results of the ablation study show that RoBERTa-CNN is robust on the SDD dataset.